A proton travels in a circular path with a radius of 4.10 mm in a magnetic field of strength 1.90 T. a) Calculate the speed of the proton. b) Determine the time required for it to make one–half of a revolution. 


1
Expert's answer
2022-01-16T13:09:32-0500

a) From the equilibrium of forces:


"\dfrac {mv^2}{R}=qvB→v=\dfrac{qBR}{m}=746.191\text{ km/s}."


b) The time is


"T_{1/2}=\dfrac{\pi R}{v}=17.3\text{ ns}."